Fine the slope of the line containing the points (1,2) and (2,6)
Answer by TimothyLamb(4379)

slope = m = dy/dx
m = (6-2)/(2-1)
m = (4)/(1)
m = 4
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---